OpenStudy (anonymous):

5m + 3n + p, -5p + 3n, and 2n - m. its been a while since ive tried this stuff and my son needs help

OpenStudy (godlovesme):

um is this simplifying ? if so all u have to do is collect like terms :3

OpenStudy (amistre64):

what is it you are trying to accomplish?

OpenStudy (anonymous):

5m + 3n + p - 5p + 3n + 2n -m Rearrange them 5m - m + 3n + 3n + 2n +p -5p = 4m +8n -4p = 4(m+2n-p)

OpenStudy (anonymous):

5m + 2n = 3 -5m +3n = 7 -------------------- 5n = 10 n = 10/5 n = 2 substitute the value of n in equation 1 5m + 2n = 3 5m + 2(2) = 3 5m + 4 = 3 5m = 3 - 4 5m = -1 m = -1/5 hence, n = 2 and m = -1/5

OpenStudy (anonymous):

Thanks guys and yes I was trying to simplify but its been quite a while